A man heard the echo of his voice by a cliff 3secs later. What is the velocity of sound in air if the cliff is 510m away.
1
Expert's answer
2019-04-10T09:43:18-0400

We can find the velocity of sound in air from the formula:


"v_{sound} = \dfrac{Total Distance}{Time} = \dfrac{2s}{t},"

here, "s" is the distance traveled by the sound to the cliff (since the sound reflects back from the cliff to the man, the total distance traveled by the sound is doubled), "t" is time.

Then, we get:


"v_{sound} = \dfrac{2s}{t} = \dfrac{2 \cdot 510 m}{3 s} = 340 \dfrac{m}{s}."

Answer:

"v_{sound} = 340 \dfrac{m}{s}."